OpenStudy (anonymous):

PLEASE HELP<3 I give medals Solve 2x + 8y = 1 by the substitution method. x = 2y

OpenStudy (bahrom7893):

GO IGBASALLOTE!

OpenStudy (lgbasallote):

2(2y) + 8y = 1 4y +8y = 1 12 y = 1 y = 1/12

OpenStudy (anonymous):

Thanks :3 You rock

OpenStudy (anonymous):

2(2y) + 8y =1 12y=1 y=1/12